# Gaussian Processes enabled model calibration in the context of deep geological disposal

### v2

@@ 1 Introduction
Deep repository material models are complex geological models that account for the mechanics of soil and rock, hydrological properties, thermal effects, and chemical interactions [1, 2, 3]. They are highly parametrized and the associated numerical analyses are computationally expensive [4, 5]. Their complexity has prevented their widespread adoption for many-query tasks, that is, simulations to explore multiple scenarios and the effects of uncertainties on the deep repository prognosis [5]. In this regard, we propose using Gaussian Processes (GPs) as surrogates for high-fidelity geological models. GPs can approximate the outputs of complex simulations with much lower computational cost, enabling efficient calibration, design and validation [6, 7, 8]. Speeding up the simulations with the use of a GP-based surrogate will not only allow for validated prognoses for the repository through the assimilation of data but also help in making informed decisions and enhancing the overall robustness and adaptability of the repository management process.
@@ 1.1 Constitutive models for deep geological repositories
Rock salt formations are one of the potential host rocks in Germany considered for secure long-term nuclear waste storage due to their distinctive mechanical and hydraulic properties [9]. Ensuring the natural integrity of geological barriers is crucial for safety, therefore reliable numerical calculations that depend on advanced material models are essential. The thermo-mechanical behavior of rock salt is typically assessed by means of laboratory tests, in particular, short-term triaxial compression tests and long-term creep tests [10, 11, 12]. In order to include all the characteristics of rock salt various constitutive models have been developed to effectively simulate the hydraulic, thermal, and mechanical behavior of rock salt. These constitutive models are able to capture for instance multiple creep phases, healing, and dependency on, e.g., stress condition, time, temperature, and humidity. Several different approaches to capturing the complex behavior of rock salt were compared and briefly described in past research projects [13, 14, 15]. Among them is the constitutive model TUBSsalt, which was developed by the Institute for Geomechanics and Geotechnics (IGG, TU Braunschweig) and presented for the first time in [16]. It has been shown in [14] and [15] that TUBSsalt, as well as other constitutive models, accurately capture the thermal and mechanical characteristics of rock salt. The application of such complex constitutive models requires expert knowledge and ideally data to infer the numerous model parameters. Therefore, this constitutive model is the focus of the presented calibration process.
@@ 1.2 Gaussian processes as efficient surrogates for computationally expensive models
Unlike much of the existing literature on surrogate modeling for complex mechanical systems, which often relies on synthetic data, our study goes beyond the implementation of a surrogate modeling approach and exploits it to enable the calibration of the TUBSsalt constitutive model using 14 years of real-world monitoring data collected in an open drift located in the northern main drift of Gorleben, Germany. The primary contribution of this paper is to demonstrate the applicability and effectiveness of a GP-based data-driven methodology in addressing a real-world problem. Specifically:
@@ 2.1 Detailed site description and problem geometry
The mechanical model considered in this work is based on the cross-section of a drift located on the north-western flank of the Gorleben salt dome, a former salt exploration mine in Germany. Monitoring data were collected and provided by the German Federal Company for Radioactive Waste Disposal (BGE mbH). Figure 1 shows a vertical geological cross section from [45] that divides the Gorleben salt dome into different homogeneous salt areas, highlighting the location of the considered drift within the salt layer known as Streifensalz (z2HS2).
Figure 1: Vertical geological cross-section of the Gorleben salt dome, showing the location of the considered drift (small red circle). The stratigraphy was obtained from the open-access report [45].
@@
Depending on crystallinity and the presence of secondary aggregates, the rock salt can be divided into homogeneous areas based on their viscous behavior, specifically the steady-state (secondary) creep rate [46]. Since the measurement location is homogeneously surrounded by a salt formation known as z2HS2, interactions between different homogeneous areas are not taken into account. Therefore, the entire numerical model can be constructed under the assumption of a uniform rock salt material.
@@
Figure 2: Cross-sectional representation of the drift area. a. Cross-section of the measurement locations from the Gorleben site, provided by the BGE mbH. b. Computational model of the drift in FLAC3D, showing the history locations used for evaluating displacements.
@@ 2.2.2 Constitutive model TUBSsalt
where $\dot{\bm{\varepsilon}}_{p}$ represents the primary creep rate tensor, $F_{p}$ the yield function, $\eta_{p}^{*}$ the current viscosity of primary creep and $\frac{\partial\sigma_{eq}}{\partial\bm{\sigma}}$ denotes the derivatives of the equivalent stress $\sigma_{eq}$ with respect to the components of the stress tensor. In case of a decrease of the stress deviator, the viscosity of primary creep $\eta_{p}^{*}$ is replaced by the viscosity for recovery creep $\eta_{p,rec}$.
@@
$\dot{\bm{\varepsilon}}_{s}=\frac{F_{s}}{\eta_{s}}\cdot\frac{\partial\sigma_{eq}}{\partial\bm{\sigma}}\cdot q_{s},$ (2.9)
where $\dot{\bm{\varepsilon}}_{s}$ represents the secondary creep rate tensor, $\eta_{s}$ is the viscosity parameter for secondary creep and $q_{s}$ is the temperature coefficient for secondary creep.
@@
$F_{t}>0:\,\dot{\bm{\varepsilon}}_{t}=\frac{F_{t}\cdot k_{t}}{\eta_{t}^{*}\cdot q_{t}}\cdot\frac{\partial Q}{\partial\bm{\sigma}},$ (2.10)
where $\dot{\bm{\varepsilon}}_{t}$ represents the tertiary creep rate tensor, $\eta_{t}^{*}$ is the current viscosity of tertiary creep, $k_{t}$ a coefficient for loading rate, $q_{t}$ a temperature coefficient and $\frac{\partial Q}{\partial\bm{\sigma}}$ the directional derivatives of the potential function for tertiary creep $Q$ with respect to the components of the stress tensor. The increase in volume or dilatancy due to microcracks is taken into account in $\frac{\partial Q}{\partial\bm{\sigma}}$.
@@
where $\sigma_{z}$ is the tensile strength, $q_{t}$ a temperature coefficient, $\eta_{v}^{*}$ the current viscosity of healing and $\frac{\partial Q_{v}}{\partial\bm{\sigma}}$ the directional derivatives of the potential function for healing $Q_{v}$ with respect to the components of the stress tensor.
@@
where $\dot{\bm{\varepsilon}}_{n}$ is the strain rate tensor of the post-failure, $k$ a coefficient for loading rate, $\eta_{n}^{*}$ the current viscosity parameter for healing and $q_{n}$ a temperature coefficient.
@@
For further details on quantities not explained in this section, such as the yield functions $F_{p}$, $F_{t}$, $F_{n}$, the potential functions for tertiary creep and healing $Q$ and $Q_{v}$, as well as derivations of all equations of TUBSsalt provided here, the reader is referred to [49]. In summary, the constitutive model TUBSsalt comprises a total of 25 material parameters, summarized in Appendix B, Table B1.
@@ 2.3 Numerical solution
The constitutive model TUBSsalt is implemented into the commercial software FLAC3D [48], which is a program for three-dimensional engineering mechanics computations. FLAC3D relies on the finite volume method to translate the continuum equations into ordinary differential equations for each element, which are then solved using an explicit central finite difference approach in time. The spatial discretization in FLAC3D is realized by meshing the continuum into hexahedral elements, with the option to use tetrahedral, wedge, and pyramid elements. It is recommended to use hexahedral elements, which are divided into tetrahedral elements to reduce the volumetric locking effect and thus achieve more accurate results. The balance of the linear momentum, equation (2.1), is iterated to an equilibrium state, which is achieved when the unbalanced mechanical force for all the gridpoints in the model is negligibly low.
@@ 5.1 High-fidelity model simulation
Monitoring data: The mechanical model described in Section 2 represents an open drift located in the northern main drift of Gorleben for which monitoring data is available. The first convergence measurement took place on October 20, 1999, one day after the excavation of the drift. This is important because it allows for the measurement of primary creep, which is most pronounced at the beginning of the excavation. On the day of the initial measurement, the horizontal distance (1-3) measured $8.94\,m$, and the vertical distance (2-4) was $6.00\,m$, as depicted in Figure 2(a). The duration of the monitoring was approximately $14.3$ years, with a significant variation of time intervals between measurements. During the first two months, measurements were taken every two to three days in order to capture the high creep rates. Subsequently, the measurement interval was increased from one week to one month, and from the end of 2002 onwards, measurements were taken approximately every six months as a constant creep rate had been achieved. Since the measuring points of the convergence distances are anchored in the rock over an anchoring length, the deformations of the cavity are also measured at these fixed points in the rock salt. Those locations are defined as history points in the numerical model in FLAC3D, demonstrated in Figure 2(b). The convergences $\mathbf{y}_{\text{moni}}^{T}(t_{i})=\left[\Delta u_{x}(t_{i}),\,\Delta u_{z}(t_{i})\right]^{T}$ are measured as a vertical and horizontal distance between two points, therefore the displacements $u_{x}$ in horizontal and $u_{z}$ in vertical direction are obtained from the displacement vector $\mathbf{u}$ for the specific nodes. The uncertainty of the convergence measurements is around ±$\pm 0.5\,mm$, while observations are on the order $\mathcal{O}(10^{2})$. Consequently, we consider the monitoring data to be noise-free.
@@
Model setup: Since only one material is considered and the cross-section of the drift is almost symmetric, only the right symmetry half of the model needs to be discretized. Dirichlet boundary conditions are applied by fixing the left, right, and bottom sides of the system in their normal directions. We model the overburden with a load of $16.774\,MPa$ applied to the top of the model, which is a Neumann boundary condition. According to the geology in Figure 1, the load results from the different heights and densities of the rock salt, cap rock, tertiary and quaternary layer. The densities were obtained from [63]. The initial stress state is derived from the overburden load and displacements for time $t=0$ are set to zero. The initial stress state is assumed to be isotropic. The slight directional dependence of the monitoring data can be attributed to the differing dimensions of the drift cross-section in the $x$ and $z$ directions. The absence of nearby geotechnical structures and the sufficient distance between the measurement site and other side drifts, or homogeneous areas allow the model to be simplified to a 2D model. The dimensions in the $x$ and $z$ direction for the model are given by $l_{x}=50\,m$ and $l_{z}=100\,m$, so that according to [64] an influence of the drift on the system boundaries can be excluded. Temperature measurements conducted in the close area of the drift show an almost constant value of $37~{}^{\circ}$. Therefore, a constant temperature is assumed over the entire system as a simplification. The mesh density of the model was optimized by performing a convergence study, ensuring that increasing the number of elements results in no significant changes in the curves of simulated horizontal and vertical convergences. Prior to the time-dependent creep analysis, the model reaches an equilibrium state twice: first, after applying the initial and boundary conditions, and second, after setting the stresses in the zones inside the cross section to zero to simulate the excavation. The equilibrium state is reached as soon as the average ratio between out-of-balance and total forces falls below $1\cdot 10^{-6}$.
@@
Sampling of parameter values in Flac3D: Following the approach in Section 3.1, a parameter set for the constitutive model TUBSsalt based on laboratory strength and creep tests has been determined in a project presented in [50] for the Gorleben site. These values serve as characteristic parameter values for Gorleben salt and also help define reasonable parameter ranges for those parameters that need calibration, see also Table B1 in Appendix B. In order to create training and testing data by means of a multitude of simulations, the parameter values of the selected TUBSsalt parameters are sampled in the given ranges in Table B1 assuming a uniform distribution. For this purpose, a SciPy Quasi-Monte Carlo generator is used to generate a Sobol sequence by creating low-discrepancy, quasi-random numbers [65, 66]. The loop for parameter sampling is implemented in the creep analysis of Flac3D, so that the saved equilibrium state of the model is recalled for every iteration. The creep analysis is performed for the same duration of the monitoring. To maintain the state of equilibrium, the average ratio between out-of-balance and total forces is limited to $5\cdot 10^{-6}$. The timestep increases over time with a maximum of $\Delta t=5.56$ hours. With the given configuration, a single simulation run of the high-fidelity model requires approximately $4$ minutes.
Processing the full-field simulation results to convergences: After carrying out the simulations to create training and testing data, the displacements derived at the history locations are converted into convergences at the time instances for which monitoring data are available. To determine the vertical convergences, the amounts of the vertical displacements $u_{z}(t_{i})$ at the top and bottom are summed up together. To determine the horizontal convergences, the amount of the horizontal displacements $u_{x}(t_{i})$ is multiplied by two to account for the symmetry of the system. The displacements of an exemplary full-field simulation at $t=14.5$ years is depicted in Figure 4. In order to compare the simulated convergences with the monitoring data, the convergences of the simulations must be set to zero at the time of the initial monitoring measurement. Therefore, the convergences from the first day after excavation of the drift can not be analysed and will be neglected. Convergences are determined at all points in time at which monitoring is available. As a result, Figure 5 shows the processed data of $200$ simulations. Figure 5a depicts the histograms obtained after sampling the input material parameters, while Figure 5b,c correspond to the vertical and horizontal convergences over time (blue lines) in comparison with the corresponding monitoring data (black squares) and the obtained probability density function (PDF) at selected time instances (red area). This figure illustrates how the uncertainty in material parameters propagates through the mechanical model implemented in FLAC3D, reflecting its nonlinearity. That is, uniformly sampled inputs result in heavily tailed outputs. The task of the GP-based surrogate is to represent this input-output relationship by learning the PDF of the convergences at different time instances as a function of the model parameters.
@@
Figure 5: Input-output simulation data at the monitoring location. a. Histograms for the input material parameters $\eta_{p}$, $E_{p}$, $\eta_{s}$, $p_{s}$, $\sigma_{0,eq,p}$, $p_{p}$, and $\sigma_{0,eq,s}$ of TUBSsalt. b. Vertical and c. horizontal convergence trajectories over time. b and c show the monitoring data (black squares) and $200$ model realizations (blue lines) from the FLAC3D simulations along with the obtained PDF at selected time instances (red area).

### v3

@@ 1 Introduction
Deep repository material models are complex geological models that account for the mechanics of soil and rock, hydrological properties, thermal effects, and chemical interactions [1, 2]. They are highly parametrized and the associated numerical analyses are computationally expensive [3, 4]. Their complexity has prevented their widespread adoption for many-query tasks, that is, simulations to explore multiple scenarios and the effects of uncertainties on the deep repository prognosis [4]. In this regard, we propose using Gaussian Processes (GPs) as surrogates for high-fidelity geological models. GPs can approximate the outputs of complex simulations with much lower computational cost, enabling efficient calibration, design and validation [5, 6, 7]. Speeding up the simulations with the use of a GP-based surrogate will not only allow for validated prognoses for the repository through the assimilation of data but also help in making informed decisions and enhancing the overall robustness and adaptability of the repository management process.
@@ 1.1 Constitutive models for deep geological repositories
Rock salt formations are one of the potential host rocks in Germany considered for secure long-term nuclear waste storage due to their distinctive mechanical and hydraulic properties [8, 9]. Ensuring the natural integrity of geological barriers is crucial for safety, therefore reliable numerical calculations that depend on advanced material models are essential [9]. The thermo-mechanical behavior of rock salt is typically assessed by means of laboratory tests, in particular, short-term triaxial compression tests and long-term creep tests [10, 11, 12]. In order to include all the characteristics of rock salt, various constitutive models have been developed to effectively simulate the hydraulic, thermal, and mechanical behavior of rock salt; refer, for example, to [13, 14, 15]. These constitutive models are able to capture for instance multiple creep phases, healing, and dependency on, e.g., stress condition, time, temperature, and humidity. Several different approaches to capturing the complex behavior of rock salt were compared and briefly described in past research projects [16, 17, 18]. Among them is the constitutive model TUBSsalt, which was developed by the Institute for Geomechanics and Geotechnics (IGG, TU Braunschweig) and presented for the first time in [19]. It has been shown in [15], [17] and [18] that TUBSsalt, as well as other constitutive models, accurately capture the thermal and mechanical characteristics of rock salt. The application of such complex constitutive models requires expert knowledge and ideally data to infer the numerous model parameters. Therefore, this constitutive model is the focus of the presented calibration process.
@@ 1.2 Gaussian processes as efficient surrogates for computationally expensive models
Unlike much of the existing literature on surrogate modeling for complex mechanical systems, which often relies on synthetic data, our study goes beyond the implementation of a surrogate modeling approach and exploits it to enable the calibration of the TUBSsalt constitutive model using $14$ yr of real-world monitoring data collected in an open drift located in the northern main drift of Gorleben, Germany. The primary contribution of this paper is to demonstrate the applicability and effectiveness of a GP-based data-driven methodology in addressing a real-world problem. Specifically:
@@ 2.1 Detailed site description and problem geometry
The mechanical model considered in this work is based on the cross-section of a drift located on the north-western flank of the Gorleben salt dome, a former salt exploration mine in Germany. Monitoring data were collected and provided by the German Federal Company for Radioactive Waste Disposal (BGE mbH).
@@
Depending on crystallinity and the presence of secondary aggregates, the rock salt can be divided into homogeneous areas based on their viscous behavior, specifically the steady-state (secondary) creep rate [48]. A vertical geological cross-section of the Gorleben salt dome and its homogeneous salt areas can be found in [49]. Since the measurement location is homogeneously surrounded by a salt formation known as Streifensalz z2HS2, interactions between different homogeneous areas are not taken into account. Therefore, the entire numerical model can be constructed under the assumption of a uniform rock salt material.
@@
Figure 1: Cross-sectional representation of the drift area. a. Cross-section of the measurement locations from the Gorleben site, provided by the BGE mbH. b. Computational model of the drift in FLAC3D, showing the history locations used for evaluating displacements and the mesh used in the numerical solution.
@@ 2.2.2 Constitutive model TUBSsalt
where $\dot{\bm{\varepsilon}}_{p}$ represents the primary creep rate tensor, $F_{p}$ the yield function of primary creep, $\eta_{p}^{*}$ the current viscosity of primary creep and $\frac{\partial\sigma_{eq}}{\partial\bm{\sigma}}$ denotes the derivatives of the equivalent stress $\sigma_{eq}$ with respect to the components of the stress tensor. In case of a decrease of the stress deviator, the viscosity of primary creep $\eta_{p}^{*}$ is replaced by the viscosity for recovery creep $\eta_{p,rec}$.
@@
$\dot{\bm{\varepsilon}}_{s}=\frac{F_{s}\cdot q_{s}}{\eta_{s}}\cdot\frac{\partial\sigma_{eq}}{\partial\bm{\sigma}},$ (2.9)
where $\dot{\bm{\varepsilon}}_{s}$ represents the secondary creep rate tensor, $F_{s}$ the yield function of secondary creep, $\eta_{s}$ is the viscosity parameter for secondary creep and $q_{s}$ is the temperature coefficient for secondary creep.
@@
$F_{t}>0:\,\dot{\bm{\varepsilon}}_{t}=\frac{F_{t}\cdot k}{\eta_{t}^{*}\cdot q_{t}}\cdot\frac{\partial Q_{t}}{\partial\bm{\sigma}},$ (2.10)
where $\dot{\bm{\varepsilon}}_{t}$ represents the tertiary creep rate tensor, $\eta_{t}^{*}$ is the current viscosity of tertiary creep, $k$ a coefficient for loading rate and stress state, $q_{t}$ a temperature coefficient and $\frac{\partial Q_{t}}{\partial\bm{\sigma}}$ the directional derivatives of the potential function for tertiary creep $Q_{t}$ with respect to the components of the stress tensor. The increase in volume or dilatancy due to microcracks is taken into account in $\frac{\partial Q_{t}}{\partial\bm{\sigma}}$.
@@
where $\sigma_{z}$ is the tensile strength, $q_{v}$ a temperature coefficient, $\eta_{v}^{*}$ the current viscosity of healing and $\frac{\partial Q_{v}}{\partial\bm{\sigma}}$ the directional derivatives of the potential function for healing $Q_{v}$ with respect to the components of the stress tensor.
@@
where $\dot{\bm{\varepsilon}}_{n}$ is the strain rate tensor of the post-failure, $\eta_{n}^{*}$ the current viscosity parameter for healing and $q_{n}$ a temperature coefficient.
@@
For further details on quantities not explained in this section, such as the yield functions $F_{p}$, $F_{t}$ and $F_{n}$, the potential functions for tertiary creep and healing $Q_{t}$ and $Q_{v}$, as well as derivations of all equations of TUBSsalt provided here, the reader is referred to [52]. In summary, the constitutive model TUBSsalt comprises a total of $25$ material parameters, summarized in Appendix B, Table B1.
@@ 2.3 Numerical solution
The constitutive model TUBSsalt is implemented into the commercial software FLAC3D [51], which is a program for three-dimensional engineering mechanics computations. FLAC3D relies on the finite difference method (FDM) to translate the continuum equations into ordinary differential equations for each element, which are then solved using an explicit central finite difference approach in time. The spatial discretization in FLAC3D is realized by meshing the continuum into hexahedral elements, with the option to use tetrahedral, wedge, and pyramid elements. It is recommended to use hexahedral elements, which are divided into tetrahedral elements to reduce the volumetric locking effect and thus achieve more accurate results. The balance of the linear momentum, equation (2.1), is iterated to an equilibrium state, which is achieved when the unbalanced mechanical force for all the gridpoints in the model is negligibly low.
@@ 5.1 High-fidelity model simulation
Monitoring data: The mechanical model described in Section 2 represents an open drift located in the northern main drift of Gorleben for which monitoring data is available. The first convergence measurement took place on October 20, 1999, one day after the excavation of the drift. This is important because it allows for the measurement of primary creep, which is most pronounced at the beginning of the excavation. On the day of the initial measurement, the horizontal distance (1-3) measured $8.94$ m, and the vertical distance (2-4) was $6.00$ m, as depicted in Figure 1(a). The duration of the monitoring was approximately $14.3$ yr, with a significant variation of time intervals between measurements. During the first two months, measurements were taken every two to three days in order to capture the high creep rates. Subsequently, the measurement interval was increased from one week to one month, and from the end of 2002 onwards, measurements were taken approximately every six months as a constant creep rate had been achieved. Since the measuring points of the convergence distances are anchored in the rock over an anchoring length, the deformations of the cavity are also measured at these fixed points in the rock salt. Those locations are defined as history points in the numerical model in FLAC3D, demonstrated in Figure 1(b). The convergences $\mathbf{y}_{\text{moni}}^{T}(t_{i})=\left[\Delta u_{x}(t_{i}),\,\Delta u_{z}(t_{i})\right]^{T}$ are measured as a vertical and horizontal distance between two points, therefore the displacements $u_{x}$ in horizontal and $u_{z}$ in vertical direction are obtained from the displacement vector $\mathbf{u}$ for the specific nodes. The uncertainty of the convergence measurements is around ±$\pm 0.5$ mm, while observations are on the order $\mathcal{O}(10^{2})$. Consequently, we consider the monitoring data to be noise-free.
@@
Model setup: Since only one material is considered and the cross-section of the drift is almost symmetric, only the right symmetry half of the model needs to be discretized. Dirichlet boundary conditions are applied by fixing the left, right, and bottom sides of the system in their normal directions. We model the overburden with a load of $16.774$ MPa applied to the top of the model, which is a Neumann boundary condition. According to the geology of the site given in [49], the load results from the different heights and densities of the rock salt, cap rock, tertiary and quaternary layer. The densities were obtained from [66]. The initial stress state is derived from the overburden load and displacements for time $t=0$ are set to zero. The initial stress state is assumed to be isotropic. The slight directional dependence of the monitoring data can be attributed to the differing dimensions of the drift cross-section in the $x$ and $z$ directions. The absence of nearby geotechnical structures and the sufficient distance between the measurement site and other side drifts, or homogeneous areas allow the model to be simplified to a 2D model. The dimensions in the $x$ and $z$ direction for the model are given by $l_{x}=50$ m and $l_{z}=100$ m, so that an influence of the drift on the system boundaries can be excluded. Temperature measurements conducted in the close area of the drift show an almost constant value of $37~{}^{\circ}$. Therefore, a constant temperature is assumed over the entire system as a simplification. The mesh density of the model was optimized by performing a convergence study, ensuring that increasing the number of elements results in no significant changes in the curves of simulated horizontal and vertical convergences. Prior to the time-dependent creep analysis, the model reaches an equilibrium state twice: first, after applying the initial and boundary conditions, and second, after setting the stresses in the zones inside the cross section to zero to simulate the excavation. The equilibrium state is reached as soon as the average ratio between out-of-balance and total forces falls below $1\cdot 10^{-6}$.
@@
Sampling of parameter values in Flac3D: Following the approach in Section 3.1, a parameter set for the constitutive model TUBSsalt based on laboratory strength and creep tests has been determined in a project presented in [53] for the Gorleben site. These values serve as characteristic parameter values for Gorleben salt and also help define reasonable parameter ranges for those parameters that need calibration, see also Table B1 in Appendix B. In order to create training and testing data by means of a multitude of simulations, the parameter values of the selected TUBSsalt parameters are sampled in the given ranges in Table B1 assuming a uniform distribution. For this purpose, a SciPy Quasi-Monte Carlo generator is used to generate a Sobol sequence by creating low-discrepancy, quasi-random numbers [67, 68]. The loop for parameter sampling is implemented in the creep analysis of Flac3D, so that the saved equilibrium state of the model is recalled for every iteration. The creep analysis is performed for the same duration of the monitoring. To maintain the state of equilibrium, the average ratio between out-of-balance and total forces is limited to $5\cdot 10^{-6}$. The timestep increases over time with a maximum of $\Delta t=5.56$ h. With the given configuration, a single simulation run of the high-fidelity model requires approximately $4$ min.
@@
Processing the full-field simulation results to convergences: After carrying out the simulations to create training and testing data, the displacements derived at the history locations are converted into convergences at the time instances for which monitoring data are available. To determine the vertical convergences, the amounts of the vertical displacements $u_{z}(t_{i})$ at the top and bottom are summed up together. To determine the horizontal convergences, the amount of the horizontal displacements $u_{x}(t_{i})$ is multiplied by two to account for the symmetry of the system. The displacements of an exemplary full-field simulation at $t=14.5$ yr is depicted in Figure 3. In order to compare the simulated convergences with the monitoring data, the convergences of the simulations must be set to zero at the time of the initial monitoring measurement. Therefore, the convergences from the first day after excavation of the drift can not be analysed and will be neglected. Convergences are determined at all points in time at which monitoring is available. As a result, Figure 4 shows the processed data of $200$ simulations. Figure 4a depicts the histograms obtained after sampling the input material parameters, while Figure 4b,c correspond to the vertical and horizontal convergences over time (blue lines) in comparison with the corresponding monitoring data (black squares) and the obtained probability density function (PDF) at selected time instances (red area). This figure illustrates how the uncertainty in material parameters propagates through the mechanical model implemented in FLAC3D, reflecting its nonlinearity. That is, uniformly sampled inputs result in heavily tailed outputs. The task of the GP-based surrogate is to represent this input-output relationship by learning the PDF of the convergences at different time instances as a function of the model parameters.
@@
Figure 4: Input-output simulation data at the monitoring location. a. Histograms for the input material parameters $\eta_{p}$, $E_{p}$, $\eta_{s}$, $p_{s}$, $\sigma_{0,eq,p}$, $p_{p}$, and $\sigma_{0,eq,s}$ of TUBSsalt. b. Vertical and c. horizontal convergence trajectories over time. b. and c. show the monitoring data (black squares) and $200$ model realizations (blue lines) from the FLAC3D simulations along with the obtained PDF at selected time instances (red area).